### v1

Abstract
Drug repositioning is an economically efficient strategy used to discover new indications for existing drugs beyond their original approvals, expanding their applicability and usage to address challenges in disease treatment. In recent years, deep-learning techniques for drug repositioning have gained much attention. While most deep learning-based research methods focus on encoding drugs and diseases by extracting feature information from neighbors in the network, they often pay little attention to the potential relationships between the features of drugs and diseases, leading to imprecise encoding of drugs and diseases. To address this, we design a dual-feature drug repositioning neural network (DFDRNN) model to achieve precise encoding of drugs and diseases. DFDRNN uses two features to represent drugs and diseases: the similarity feature and the association feature. The model incorporates a self-attention mechanism to design two dual-feature extraction modules for achieving precisely encoding of drugs and diseases: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. The IntraDDFE module extracts features from a single domain (drug or disease domain), while the InterDDFE module extracts features from the mixed domain (drug and disease domain). In particular, the feature is changed by InterDDFE, ensuring a precise encoding of drugs and diseases. Finally, a cross-dual-domain decoder is designed to predict drug-disease associations in both the drug and disease domains. Compared to six state-of-the-art methods, DFDRNN outperforms others on four benchmark datasets, with an average AUROC of 0.946 and an average AUPR of 0.597.
keywords:
Drug repositioning, neural network, dual-feature, association feature, cross-dual-domain decoder
1 Introduction
In today’s Medical field, drug development is complex and time-consuming. Developing a new drug involves three stages: the discovery stage, the preclinical stage, and the clinical stage, which typically takes 10-15 years, with an average cost of up to \$2.8 billion fisher2006alterations ; vijayan2022enhancing . Although scientists continue to discover new drug molecules through unremitting efforts and innovation, the traditional drug development model often faces many risks, including the termination of drug development due to drug efficacy, drug safety, or commercial concerns. Remarkably, a significant proportion (80-90%) of these projects fail in clinical trials vijayan2022enhancing ; dimasi2010trends . Therefore, finding a more efficient and precise method for drug development has become critically important.
Drug repositioning, a strategy that garnered significant attention, is proposed to explore alternative indications and potential side effects for already-licensed medications dudley2011exploiting ; wang2024landscape . Compared with traditional drug research and development methods, this approach offers several advantages, such as shorter time, lower cost, and reduced risk. Computational drug repositioning, a favored method among researchers, can significantly reduce the time and energy required to identify candidate drugs. Computational drug repositioning methods are divided into four categories cai2023machine ; luo2021biomedical : network diffusion-based methods, machine learning-based methods, matrix factorization-based methods, and deep learning-based methods.
Network diffusion-based methods classify the drug reposition-ing problem as a link prediction issue within graph neural networks kumar2020link . These methods focus on building heterogeneous networks containing multiple types of biological entities and linking drugs to diseases through message propagation on paths bridging different networks. A commonly used method is a random walk and its variants luo2021biomedical ; lagcn . For instance, Martinez et al. martinez2015drugnet proposed a network-based drug repositioning prioritization method called DrugNet. It integrates information about diseases, drugs, and proteins to simultaneously prioritize drug-disease and disease-drug associations. Another example is Luo et al.’s method BIRW_Cdataset , which effectively predicted potential associations between drugs and diseases by constructing an integrated similarity network between drugs and diseases and combining it with a bidirectional random walk (BiRW).
Machine learning-based methods classify drug repositioning as a binary classification problem. These methods typically use traditional machine learning techniques for drug-disease association prediction, including random forest (RF), support vector machine (SVM), and decision trees cai2023machine . For example, Peng et al. screening designed a novel DTI screening framework called PUDTI. PUDTI utilizes SVM to infer candidate targets for drug repositioning. Shi et al. shi2019lasso proposed a method for predicting drug-target interactions based on Lasso dimensionality reduction and a random forest classifier. Zhao et al. HINGRL proposed HINGRL, which uses a random forest classifier in the final decoding step to predict drug-disease association. In addition, there are decision trees gradient , linear regression, and logistic regression you2019predicting ; elisei2019recursive .
Matrix decomposition-based methods categorize the drug repositioning problem as a recommendation system issue. These methods employ matrix factorization techniques to reduce dimensionality and extract key features, thus better capturing the associations between drugs and diseases. Some standard matrix factorization techniques include singular value decomposition (SVD) and non-negative matrix factorization (NMF). For example, Yang et al. MSBMF proposed a Multiple Similarity Bilinear Matrix Factorization (MSBMF) method, where the similarity matrices of drugs and diseases are concatenated and the drug-disease association matrix is reshaped through matrix decomposition. Luo et al. luo2018SVT employed a fast Singular Value Thresholding (SVT) algorithm to propose a drug repositioning recommendation system (DRRS). Additionally, Sadegh et al. sadeghi2022NMF-DR and Xuan et al. xuan2019dd combined non-negative matrix factorization to predict drug-disease associations.
In recent years, methods based on deep learning have become a hot research topic for drug repositioning. They can effectively capture high-level hidden representations from various data sources and demonstrate state-of-the-art performance. Unlike traditional machine learning methods, deep learning approaches can learn more abstract information by building deep architectures without manual feature selection and tuning yu2022deep . Generally, methods based on deep learning for drug repositioning often integrate graph neural networks such as Graph Convolutional Networks (GCN) and Graph Attention Networks (GAT) to construct models. For example, Yu et al. lagcn integrated a drug-disease heterogeneous network with a GCN network to build a multi-layer connected end-to-end network model called LAGCN. They also incorporated layer attention mechanisms for drug-disease association prediction. Gu et al. gu2022REDDA designed a biological network of five biological entities and equipped it with three attention mechanisms to predict drug-disease associations. Wang et al. wang2023MSGCA designed a multi-similarity graph convolutional autoencoder(MSGCA). Sun et al. sun2022PSGCN predicted new drug-disease associations by constructing partner-specific graphs for drug-disease pairs and combining them with the GCN. By incorporating the GCN, Xuan et al. xuan2022heterogeneous constructed multiple drug-disease heterogeneous networks based on different drug similarities and encoded drugs and diseases at Multi-scale neighbors. Cai et al. cai2021DRHGCN designed intra-domain and inter-domain feature extraction modules through the GCN and the bilinear aggregator, proposing DRHGCN. Meng et al. DRAGNN utilized a GAT for encoding drugs and diseases and introduced a weighted local information augmented graph neural network model named DRAGNN. Most current deep learning-based methods for drug repositioning typically utilize the similarity between neighbors and aggregate neighbor features for training, achieving considerable results. However, many end-to-end deep learning models do not involve purposeful encoding. They delegate the task to the optimizer, relying solely on the coordinated training of the encoding and decoding processes to complete association prediction. This may lead to the introduction of a large amount of noisy information during model training, resulting in imprecise encoding of drugs and diseases and affecting the prediction results.
This paper proposes a dual-feature drug repositioning neural network (DFDRNN) model to address the above issue. The DFDRNN constructs two features for drugs and diseases: the similarity feature and the association feature. It utilizes the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module to extract drug and disease features and achieve precise encoding. A self-attention mechanism (SAM) is employed in these two modules to capture the complex relationships between adjacent nodes. Finally, the decoder adopts a cross-dual-domain decoder to achieve precise prediction of drug-disease associations. Compared to GCN-based methods, DFDRNN employs a SAM to dynamically adjust the attention level of each input element to neighbor information, providing fine-grained weight allocation. This enables efficient aggregation of neighbors and enhances the model’s predictive power. We compare DFDRNN with six state-of-the-art methods on four benchmark datasets. Experimental results show that our model outperforms other models. This article makes the following contributions:
(1) We propose a dual-feature method to represent drugs and diseases. The proposed approach involves two features: the similarity feature and the association feature. These two features are independent of each other in the network, while they can be extracted simultaneously to represent drugs and diseases in each network layer.
(2) Two dual-feature extraction modules are proposed: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. In particular, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature. The proposed two modules effectively capture the features of drugs and diseases, ensuring a precise encoding of drugs and diseases.
(3) A cross-dual-domain decoder is proposed, implementing a cross-computation between two encoded features to obtain decoding results for the drug and disease domains. The fusion of the two decoding results improves the accuracy of association prediction.
The remainder of this paper is organized as follows. Section 2 describes the datasets and methods, detailing the implementation process of DFDRNN; Section 3 clarifies the experimental design and analyzes the experimental results; and Section 4 concludes our work.
2 Materials and Methods
2.1 Datasets
We conducted experiments on four benchmark datasets. The first dataset, Fdataset Fdataset , includes 593 drugs from DrugBank knox2024drugbank and 313 diseases listed in the Online Mendelian Inheritance in Man (OMIM) hamosh2005OMIM database, comprising 1933 drug-disease associations. The second dataset, Cdataset BIRW_Cdataset , consists of 663 drugs registered in DrugBank, 409 diseases from the OMIM database, and 2532 known drug-disease associations. The third dataset, Ldataset lagcn , contains 18,416 drug-disease associations between 269 drugs and 598 diseases sourced from the Comparative Toxicogenomics Database(CTD) davis2023CTD . The fourth dataset, LRSSL liang2017lrssl , comprises 763 drugs, 681 diseases, and 3051 validated drug-disease associations. Table 1 summarizes the four datasets’ statistics.
This study determines drug similarity based on chemical structures obtained from DrugBank using standardized SMILES notation weininger1988smiles . The Tanimoto score of the 2D chemical fingerprints of two drugs is then calculated using the chemistry development kit steinbeck2003chemistry . Disease similarity, on the other hand, is based on disease phenotypes and is computed using MimMiner van2006text . MimMiner gauges disease similarity by analyzing the similarity of MeSH terms in medical descriptions of diseases from the OMIM database lipscomb2000medical .
Table 1: The statistics of the four datasets used in this study
[…]
Cdataset 663 409 2532
Ldataset 269 598 18416
LRSSL 763 681 3051
[…]
2.2 Model structure
Our research is focused on predicting potential associations between drugs and diseases. To this end, we proposed a dual-feature drug repositioning neural network (DFDRNN) model. The encoder of DFDRNN encompasses two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module. While the IntraDDFE module extracts features among drugs (or diseases), the InterDDFE module extracts features between drugs and diseases and changes the feature. A self-attention (SAM) model is employed when aggregating information to capture the complex relationship among adjacent nodes. The decoder of DFDRNN receives the results from the encoder and implements cross-dual-domain decoding, including both the drug and disease domains. Figure 1 illustrates the diagram of the proposed DFDRNN model.
Figure 1: DFDRNN generates the similarity feature and the association feature for drugs and diseases based on established similarity networks and association networks. The encoding process utilizes a multi-layer connected network architecture, with each layer performing IntraDDFE and InterDDFE based on self-attention. A layer attention mechanism is also introduced to compute the final similarity and association features. Finally, the decoder calculates drug-disease association scores by decoding separately for the drug and disease domains.
2.2.1 Three networks
To illustrate the relationships among biological entities (drugs and diseases) and describe their feature, we establish three networks: the drug-drug similarity network, a weighted complete graph with a node-set representing the drugs, and the weight of each edge representing the degree of similarity of the two drug incident with the edge; the disease-disease similarity network, a weighted complete graph with node set representing the diseases and the weight of each edge representing the degree of similarity of the two diseases incident with the edge; and the drug-disease association network, a bipartite network with two parts representing the sets of drugs and diseases, respectively and each edge representing the associativity between the drug and diseases incident with the edge. We use three matrices to represent these three networks.
Specifically, suppose that there are $n$ drugs (denoted by $r_{1},r_{2},\ldots,r_{n}$) and $m$ diseases (denoted by $d_{1},d_{2},\ldots,d_{m}$). We use $S^{r}$ (or $S^{d}$) to denote the adjacency matrix of a drug-drug similarity network (or disease-disease similarity network), which is a $n\times n$ ($m\times m$) real matrix such that the entry $s_{ij}\in[0,1]$ denotes the degree of similarity of drug $i$ and drug $j$ (or disease $i$ and disease $j$), where $i,j\in\{1,2,\ldots,n\}$ (or $i,j\in\{1,2,\ldots,m\}$). For a drug-disease association network, we use a $n\times m$ binary matrix, denoted by $A$, to represent it, where the entry $a_{ij}=1$ if drug $i$ is associated with disease $j$ and $a_{ij}=0$ if drug $i$ is not associated with disease $j$ or the associativity is yet to be observed, for $i=1,2,\ldots,n$ and $j=1,2,\ldots,m$.
To reduce the impact of data noise, we aggregate only the top-$t$ neighbors for each drug in $S^{r}$ and each disease in $S^{d}$, and then two new matrices $R\in\mathbb{R}^{n\times n}$ and $D\in\mathbb{R}^{m\times m}$ are generated to replace $S^{r}$ and $S^{d}$, respectively, where the entries $r_{ij}$ ($i,j\in\{1,2,\ldots,n\}$) of $R$ and $d_{ij}$ ($i,j\in\{1,2,\ldots,m\}$) of $D$ are defined as follows. For a drug $r_{i}$ (and a disease $d_{i}$), we denote by $N_{k}(r_{i})$ (and $N_{k}(d_{i})$) the set of the $t$-nearest neighbors of $r_{i}$ in $S^{r}$ (and $d_{i}$ in $S^{d}$), i.e., the weight of edge $r_{i}x$ (and edge $d_{i}y$) for every $x\in N_{k}(r_{i})$ (and every $y\in N_{k}(d_{i})$) is top-$t$ among all neighbors of $r_{i}$ (and $d_{i}$), and let $\tilde{N_{k}}(r_{i})=N_{k}(r_{i})\cup\{r_{i}\}$ (and $\tilde{N_{k}}(d_{i})=N_{k}(d_{i})\cup\{d_{i}\}$) be the extended $t$-nearest neighbors of $r_{i}$ (and $d_{i}$).
$\begin{matrix}R_{ij}=\begin{cases}1,&\mbox{if }r_{j}\in{\tilde{N_{k}}(r_{i})}\\ 0,&\text{otherwise}\end{cases}\quad&D_{ij}=\begin{cases}1,&\mbox{if }d_{j}\in{\tilde{N_{k}}(d_{i})}\\ 0,&\text{otherwise}\end{cases}\end{matrix}$ (1)
Finally, to encode drugs and diseases simultaneously and distinguish IntraDDFE from InterDDFE, we construct two heterogeneous networks $E_{s}$ and $E_{a}$ based on the above matrices $R,D$, and $A$. $E_{s}$ is used for IntraDDFE, while $E_{a}$ is used for InterDDFE, defined as:
$\ \qquad\qquad E_{s}=\begin{bmatrix}R&0\\ 0&D\end{bmatrix}\qquad E_{a}=\begin{bmatrix}0&A\\ A^{T}&0\end{bmatrix}$ (2)
2.2.2 Self-attention mechanism
The self-attention mechanism (SAM) can be described as a process associated with three vectors, i.e., one query vector $Q_{i}$, some key vectors $K_{j}$, and some value vector $V_{j}$, which maps $Q$ and a group of $K$-$V$ pairs to a new output vector. The output of the attention is a weighted sum of $V$ vectors, where the weight for each $V$ is computed by taking the dot product of the $Q$ with the corresponding $K$ vaswani2017attention . In our IntraDDFE and InterDDFE modules, we use the multi-head SAM to aggregate neighbor information for drugs and diseases. We use the notation sam to represent the multi-head SAM mapping that involves two parameters $E$ and $H^{in}$, where $E\in\{E_{s},E_{a}\}$ and $H^{in}\in\mathbb{R}^{(n+m)\times k}$ is an input feature matrix with embedding dimension $k$. The specific implementation of sam can be described as follows:
$Q=H^{in}w^{Q},\qquad K=H^{in}w^{K},\qquad V=H^{in}w^{V}$ (3)
where $w^{Q},w^{K},w^{V}\in\mathbb{R}^{\mathit{k}\times k}$ are three trainable $k\times k$ real matrices. Clearly, $Q,K,$ and $V$ are $(n+m)\times k$ real matrices.
Each $Q$, $K$, and $V$ is partitioned into $p$ parts (called heads), denoted by $Q^{i}$,$K^{i}$, and $V^{i}$, $i=1,2,\ldots,p$, and the dimension of each part (i.e., the number of columns) is $k^{\prime}=k/p$. We use $Q^{q}_{i}$, $K^{q}_{i}$, and $V^{q}_{i}$ to denote the feature of node $i$ in $Q^{q}$, $K^{q}$, and $V^{q}$ (i.e., the row vector corresponding to $i$), respectively. Let $X^{q}$ be the $q$-th head of $X\in\{Q,K,V\}$. Then, the weight of aggregating neighbor information is calculated based on $Q^{q}$ and $K^{q}$. Specifically, for a node $i$ (drug or disease), the weight $\beta^{r}_{ij}$ of $i$ aggregating its neighbor $j$ is defined as
$\beta^{q}_{ij}=\frac{exp(\frac{Q^{q}_{i}\bigodot K^{q}_{j}}{\sqrt{k^{\prime}}})}{\sum_{z\in N(i)}exp(\frac{Q^{q}_{i}\bigodot K^{q}_{z})}{\sqrt{k^{\prime}}})}$ (4)
[…]
$H_{\texttt{sam}}^{i}=\texttt{concat}(\sum_{j\in N(i)}\beta^{1}_{ij}V^{1}_{j},\sum_{j\in N(i)}\beta^{2}_{ij}V^{2}_{j},\ldots,\sum_{j\in N(i)}\beta^{p}_{ij}V^{p}_{j})$ (5)
Moreover, let $H_{\texttt{sam}}$ represent the feature matrix obtained by SAM, $H_{r}\in\mathbb{R}^{n\times k}$ be the drug feature obtained by SAM, and $H_{d}\in\mathbb{R}^{m\times k}$ is the disease feature obtained by SAM. Then, we have
$H_{r}=\begin{bmatrix}H^{r_{1}}_{\texttt{sam}}\\ H^{r_{2}}_{\texttt{sam}}\\ \dots\\ H^{r_{m}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{d}=\begin{bmatrix}H^{d_{1}}_{\texttt{sam}}\\ H^{d_{2}}_{\texttt{sam}}\\ \dots\\ H^{d_{n}}_{\texttt{sam}}\end{bmatrix},\hskip 8.5359ptH_{\texttt{sam}}=\begin{bmatrix}H_{r}\\ H_{d}\end{bmatrix}$ (6)
A fully connected layer is applied after the SAM function to extract further features of drugs and diseases. We use the notation samf to denote the whole process, i.e., samf is the combination of sam with a fully connected (FC) layer. Let $w_{r}\in\mathbb{R}^{k\times k}$ and $w_{d}\in\mathbb{R}^{k\times k}$ be trainable parameter matrices, $b_{r}$ and $b_{d}$ be biases, and $\sigma$ be a LeakyRELU maas2013leakyrelu activation function. Then,
$\texttt{samf}(E,H^{in})$ $=\texttt{FC}(\texttt{sam}(E,H^{in}))$ (7)
$=\begin{bmatrix}\sigma(H_{r}w_{r}+b_{r})\\ \sigma(H_{d}w_{d}+b_{d})\end{bmatrix}$
2.2.3 Encoder
The DFDRNN model is a multi-layer connected network architecture, where each layer involves two types of dual-feature extraction modules: IntraDDFE module based on network $E_{s}$ and InterDDFE module based on network $E_{a}$. First, the initial features of drugs and diseases are generated in two features, including the similarity feature $H^{init}_{s}$ and the association feature $H^{init}_{a}$, described as follows.
$\ \qquad\qquad H^{init}_{s}=\begin{bmatrix}S^{r}&0\\ 0&S^{d}\end{bmatrix}\qquad H^{init}_{a}=\begin{bmatrix}0&A\\ A^{T}&0\end{bmatrix}$ (8)
Next, DFDRNN projects the initial features to a $k$-dimensional space through a linear layer. The similarity feature $H_{s}^{0}$ and the association feature $H_{a}^{0}$ in the 00-layer are defined as follows:
$H^{0}_{s}=H^{init}_{s}\times M\quad,\quad H^{0}_{a}=H^{init}_{a}\times M$ (9)
where $M\in\mathbb{R}^{(n+m)\times k}$ is trainable parameter matrix.
Now, we consider the process of extracting features in the $\ell$-th layer for $\ell\geq 1$. In the $\ell$-th layer, we denote by $f^{\ell}$ the IntraDDFE function in this layer, $H^{\ell}_{s}$ and $H^{\ell}_{a}$ are the input similarity feature and the input association feature, respectively, $\hat{H}^{\ell+1}_{s}$ and $\hat{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by IntraDDFE, and $\texttt{samf}^{\ell}_{f}$ the aggregation function of $\ell$-th layer for IntraDDFE. Then,
$(\hat{H}^{\ell+1}_{s},\hat{H}^{\ell+1}_{a})$ $=f^{\ell}(E_{s},H^{\ell}_{s},H^{\ell}_{a})$ (10)
$=(\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{s}),\texttt{samf}^{\ell}_{f}(E_{s},H^{\ell}_{a}))$
For the InterDDFE module, we use $g^{\ell}$ to denote the InterDDFE function in the $\ell$-th layer. Similarly, $\tilde{H}^{\ell+1}_{s}$ and $\tilde{H}^{\ell+1}_{a}$ represent the obtained similarity feature and association feature by InterDDFE, respectively, and $\texttt{samf}^{\ell}_{g}$ represents the aggregation function. Then,
$(\tilde{H}^{\ell+1}_{a},\tilde{H}^{\ell+1}_{s})$ $={g^{\ell}(E_{a},H^{\ell}_{s},H^{\ell}_{a})}$ (11)
${=(\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{s}),\texttt{samf}^{\ell}_{g}(E_{a},H^{\ell}_{a}))}$
It should be noted that compared with the IntraDDFE module, the InterDDFE module changes the features: the similarity feature $H^{\ell}_{s}$ is changed into the association feature $\tilde{H}^{\ell+1}_{a}$ and the association feature $H^{\ell}_{a}$ is changed into the similarity feature $\tilde{H}^{\ell+1}_{s}$.
Finally, for each module, the similarity (and association) features (including one input feature and two output features corresponding to the two modules) in the $\ell$-layer are added into an input similarity (and association) feature of the $(\ell+1)$-layer, i.e.,
$H^{\ell+1}_{s}=\hat{H}^{\ell+1}_{s}+\tilde{H}^{\ell+1}_{s}+H^{\ell}_{s}$ (12)
$H^{\ell+1}_{a}=\hat{H}^{\ell+1}_{a}+\tilde{H}^{\ell+1}_{a}+H^{\ell}_{a}$ (13)
As the depth of network layers increases, the valuable information for predicting drug-disease association extracted by the DFDRNN model decreases. Therefore, we introduce a layer attention mechanism to allocate different weights to the results of different layers. We use $L$ to represent the total number of layers in DFDRNN. The weight of the $\ell$-th layer, denoted by $\beta^{\ell}$, is initialized by $1/L$ and automatically learned by the model. Then, the final encoding of the similarity feature and the association feature $H^{r}_{s}$, $H^{r}_{a}$, $H^{d}_{s}$, and $H^{d}_{a}$ of drugs and diseases are obtained by implementing the layer attention mechanism as follows.
$\begin{bmatrix}H^{r}_{s}\\ H^{d}_{s}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{s}),\qquad\begin{bmatrix}H^{r}_{a}\\ H^{d}_{a}\end{bmatrix}=\sum_{\ell=1}^{L}(\beta^{\ell}\times H^{\ell}_{a})$ (14)
2.2.4 Decoder
A cross-dual-domain decoder is utilized to predict drug-disease association. We perform cross computation between two features, using $s_{r}\in\mathbb{R}^{n\times m}$ and $s_{d}\in\mathbb{R}^{n\times m}$ to represent the prediction values for the drug domain and the disease domain, respectively. Let $s\in\mathbb{R}^{n\times m}$ denote the final score matrix of drug-disease associations, where $s_{ij}$ represents the prediction score between drug $r_{i}$ and disease $d_{j}$. That is,
${s_{r}}$ $=\texttt{sigmod}({H^{r}_{a}\times(H^{d}_{s})^{T}})$ (15)
${s_{d}}$ $=\texttt{sigmod}(H^{d}_{a}\times(H^{r}_{s})^{T})$ (16)
${s}$ $={(s_{r}+(s_{d})^{T})/2.0}$ (17)
Algorithm 1 gives a detailed explanation of predicting novel drug-disease associations using DFDRNN.
Algorithm 1 DFDRNN
1:Input: Two heterogeneous network ($E_{s}$ and $E_{a}$); initial the similarity feature and the association feature (${H}^{init}_{s}$ and ${H}^{init}_{a}$); number of layers $L$; maximum training epochs $C$.
2:Output: Drug-disease association score matrix $s$.
3:Initialize the weight matrix ($w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$).
4:for $c=1$ to $C$ do
5:     Compute the similarity feature $H^{0}_{s}$ and the association feature $H^{0}_{a}$ with Equation (9).
6:     for $i=1$ to $L$ do
7:         Compute IntraDDFE $\hat{H}^{\ell}_{s},\hat{H}^{\ell}_{a}$ with Equation (10).
8:         Compute InterDDFE $\tilde{H}^{\ell}_{s},\tilde{H}^{\ell}_{a}$ with Equation (11).
9:         Fuse the same feature and the output of the previous layer yields $H^{\ell}_{s}$ and $H^{\ell}_{a}$ with Equation (12) (13)
10:     end for
11:     Utilizing layer attention mechanism to compute the final embedding with Equation (14)
12:     Obtain Computer the prediction score matrix $s$ with Equation (15)(16)(17)
13:     Update the parameters $w^{Q},w^{K},w^{V},w_{r},w_{b},\ldots$ with Equation (18).
14:     $c\leftarrow c+1$.
15:end for
2.3 Optimization
Predicting the drug-disease association can be viewed as a binary classification task. The objective is to determine whether a given drug is associated with a specific disease. Given the serious imbalance between the number of positive samples (drug-disease associations) and negative samples (drug-disease non-associations), we employ binary cross-entropy as the loss function. We use $y^{+}$ and $y^{-}$ to represent the positive and negative labels in the dataset, respectively. We use $\lambda=\frac{|y^{-}|}{|y^{+}|}$ to reduce the impact of this kind of data imbalance and denote by $(i,j)$ a drug($i$)-disease($j$) pair in $y^{+}$ or $y^{-}$, where $|y^{-}|$ and $|y^{+}|$ represent the cardinalities of $y^{+}$ or $y^{-}$, respectively.
$loss=-\frac{1}{n+m}(\lambda\times\sum_{(i,j)\in y^{+}}\log(s_{i,j})+\sum_{(i,j)\in y^{-}}\log(1-s_{i,j}))$ (18)
We optimize the DFDRNN model using the Adam kingma2014adam optimizer and initialize the parameters’ weights with Xavier initialization glorot2010understanding . We randomly drop some edges and apply regular dropout srivastava2014dropout to improve the model’s generalization ability.
3 Experiments and Results
3.1 Baseline methods
To assess DFDRNN’s performance, we compared it with six state-of-the-art drug-disease association prediction models across the above four datasets. These models are:
$\bullet$ LAGCN lagcn : This method utilizes multi-layer GCNs for neighbor information propagation on heterogeneous networks and introduces the layer attention mechanism.
$\bullet$ DRHGCN cai2021DRHGCN : This model uses GCN and bilinear aggregators to design inter-domain and intra-domain embeddings for predicting drug-disease associations.
$\bullet$ DRWBNCF meng2022DRWBNCF : This model focuses on the interaction information between neighbors and uses generalized matrix factorization collaborative filtering to predict drug-disease associations.
$\bullet$ NCH-DDA zhang2024NCH-DDA : This method involves the design of a single-neighborhood feature extraction module and a multi-neighborhood feature extraction module to extract features for drugs and diseases. Additionally, it introduces neighborhood contrastive learning for drug-disease association prediction.
$\bullet$ HDGAT huang2024HDGAT : This approach combines graph convolutional neural networks with bidirectional extended short-term memory networks and integrates layer attention mechanisms and residual connections.
$\bullet$ DRGBCN tang2023DRGBCN : This method utilizes layer attention graph convolutional networks to encode drugs and diseases and employs bilinear attention network modules to enhance their intricate relationship.
3.2 Experimental parameter configuration
The model’s parameter values are as follows: DFDRNN’s feature embedding dimension is 128; the number of network layers is 3; the number of attention heads is p = 2; the dropout rate is 0.4; the edge dropout rate is 0.5 on the Ldataset and 0.2 on other datasets; the learning rate is 0.008; and the maximum training epoch is 800. Additionally, we set top-$t$ to 7 based on parameter sensitivity experiments.
3.3 Evaluation metrics
Table 2: The statistics of the four datasets used in this study
Actual Predicted Positive Negative
Positive TP FN
Negative FP TN
We categorize the prediction results into four sets. As shown in Table 2, TP represents the number of true positives correctly classified as positive, FN represents the number of false negatives incorrectly classified as negative, FP represents the number of false positives incorrectly classified as positive, and TN represents the number of true negatives correctly classified as negative. We then calculate True Positive Rate (TPR or Recall), False Positive Rate (FPR), and precision as follows:
[…]
$\text{FPR}=\frac{FP}{FP+TN}$ (20)
$\text{Precision}=\frac{TP}{TP+FP}$ (21)
Calculating the FPR, TPR, and precision at various thresholds, we construct the receiver operating characteristic (ROC) curve, with FPR on the X-axis and TPR on the Y-axis. Additionally, we generate the precision-recall (PR) curve, with recall on the X-axis and precision on the Y-axis. The evaluation metrics used for DFDRNN are the area under the ROC curve (AUROC) and the area under the PR curve (AUPR).
3.4 Performance of DFDRNN in the cross-validation
We employ the 10-fold cross-validation method to assess the model’s performance. During this process, the dataset is randomly partitioned into ten mutually exclusive subsets, each containing 10% known drug-disease association pairs and 10% unknown pairs. One subset serves as the test set, while the remaining nine subsets are used to train the model. This process is repeated ten times (i.e., 10-fold cross-validation) with different random seeds, and the average and standard deviation of the results are calculated to compare the model’s performance. The experimental results, including AUROC and AUPR values, are presented in Table 3. The ROC and PR curves can also be found in Supplementary Figures S1-S4.
In Table 3, it is evident that DFDRNN outperformed all other models by achieving the highest average AUROC of 0.946 and the highest average AUPR of 0.597 across the four datasets. This performance surpassed the second-ranked NCH-DDA model by 0.75% and 3.1% in AUROC and AUPR, respectively. DFDRNN demonstrated the best AUROC and AUPR on all four datasets, showcasing its predictive solid capability for potential drug-disease associations based on known associations. It’s worth mentioning that NCH-DDA performed exceptionally well, coming in second only to DFDRNN, which may be attributed to its feature extraction from different neighborhoods and the fusion of features using contrastive learning loss.
Table 3: The AUROCs and AUPRs of DFDRNN and the other comparative models under 10-times 10-fold cross-validation.
[…]
AUROC LAGCN 0.881±0.002 0.911±0.002 0.864±0.001 0.935±0.001 0.898
DRHGCN 0.949±0.001 0.965±0.001 0.864±0.001 0.960±0.001 0.935
DRWBNCF 0.924±0.002 0.942±0.001 0.821±0.001 0.935±0.001 0.906
NCH-DDA 0.954±0.002 0.959±0.002 0.879±0.001 0.963±0.002 0.939
HDGAT 0.917±0.002 0.939±0.001 0.874±0.001 0.936±0.001 0.916
DRGBCN 0.931±0.003 0.946±0.001 0.824±0.002 0.943±0.001 0.911
DFDRNN 0.960±0.001 0.973±0.001 0.885±0.001 0.965±0.001 0.946
AUPR LAGCN 0.197±0.005 0.307±0.008 0.522±0.002 0.254±0.002 0.320
DRHGCN 0.531±0.003 0.619±0.003 0.523±0.001 0.408±0.002 0.520
DRWBNCF 0.493±0.001 0.569±0.004 0.413±0.003 0.349±0.005 0.456
NCH-DDA 0.623±0.006 0.675±0.003 0.552±0.002 0.466±0.002 0.579
HDGAT 0.505±0.005 0.566±0.008 0.546±0.003 0.403±0.005 0.505
DRGBCN 0.408±0.007 0.451±0.006 0.434±0.005 0.262±0.006 0.388
DFDRNN 0.623±0.003 0.688±0.002 0.573±0.001 0.502±0.002 0.597
[…]
{tablenotes}
The best results in each row are in bold faces and the second best results are underlined.
3.5 Effect of the number of nearest neighbors
Figure 2: (a)Line plot of AUROC as a function of top-$t$ on four benchmark datasets; (b)Line plot of AUPR as a function of top-$t$ on four benchmark datasets; (c)Line plot of the average AUROC as a function of top-$t$ on four datasets; (d)Line plot of the average AUPR as a function of top-$t$ on four datasets.
In the context of DFDRNN, the value of top-$t$ determines the number of neighbors features aggregated during IntraDDFE, which influences the model’s performance. A small value of $t$ limits the neighbor information obtained, thereby affecting the model’s performance. On the other hand, as $t$ increases, more neighbors are obtained, but this also introduces noise information, leading to the propagation of unnecessary data. The performance of DFDRNN under 10-fold cross-validation varies with $t$ on the four datasets, as demonstrated in Figure 2 for $t\in\{1,2,3,\ldots,20\}$.
Based on the results shown in Figure 2 (a) and (b), DFDRNN performs the poorest across the four datasets when $t=1$. However, as the $t$ value increases, DFDRNN consistently maintains excellent performance, showcasing its remarkable stability, which can be credited to its robust self-attention aggregation capability. Furthermore, the changes in average AUROC (Figure 2 (c)) and average AUPR (Figure 2 (d)) indicate that when $t=7$, DFDRNN achieves the highest average AUROC of 0.9452 and the highest average AUPR of 0.5973 across four datasets. For detailed numerical values, please refer to Supplementary Tables S1 and S2.
3.6 Ablation analysis
3.6.1 Encoder
The encoder of DFDRNN incorporates SAM to encode drugs (or diseases) into the similarity feature (SF) and the association feature (AF). We introduce two dual-feature extraction modules, IntraDDFE and InterDDFE, to precisely encode drugs and diseases, with InterDDFE changing the feature (CF) of drugs (or diseases). We developed various model variants to underscore these elements’ importance in the DFDRNN encoder and evaluated their performance.
  • DFDRNN-SF: Remove the association feature and encode drugs and diseases solely as the similarity feature.
  • DFDRNN-AF: Remove the similarity feature and encode drugs and diseases solely as the association feature.
  • DFDRNN-GCN: SAM was replaced by a GCN to observe the impact on the model’s performance.
  • DFDRNN-noCF: Remove the change of features in InterDDFE, meaning that the similarity feature and the association feature will not be changed into each other after InterDDFE.
Table 4 presents the AUROC and AUPR values of DFDRNN and its four variants under 10-fold cross-validation. By comparing the two variants of DFDRNN, DFDRNN-SF and DFDRNN-AF, we observe that DFDRNN-SF performs better than DFDRNN-AF on the Fdataset, Cdataset, and LRSSL, indicating the contribution of the similarity feature is greater than that of the association feature on these three datasets. However, on the Ldataset, the contribution of the association feature is greater than that of the similarity feature. Compared to DFDRNN, simultaneously encoding both features can significantly improve the model’s performance. Compared to DFDRNN, DFDRNN-GCN demonstrates a noticeable decrease in AUROC values, highlighting the strong performance of SAM in extracting neighbor information. Furthermore, compared to DFDRNN, DFDRNN-noCF shows a significant decrease in AUPR on the Fdataset, Cdataset, and LRSSL datasets, and performs the worst on the Ldataset, indicating the necessity of changing features in the InterDDFE module.
Table 4: The AUROCs and AUPRs of DFDRNN and its four variants under 10-fold cross-validation
[…]
AUROC AUPR AUROC AUPR AUROC AUPR AUROC AUPR
DFDRNN-SF 0.958 0.612 0.972 0.690 0.873 0.538 0.964 0.464
DFDRNN-AF 0.942 0.606 0.962 0.678 0.884 0.570 0.950 0.504
DFDRNN-GCN 0.950 0.620 0.964 0.671 0.874 0.551 0.956 0.483
DFDRNN-noCF 0.959 0.611 0.970 0.676 0.868 0.523 0.963 0.462
DFDRNN 0.960 0.622 0.973 0.689 0.885 0.572 0.965 0.504
[…]
3.6.2 Decoder
Figure 3: non-cross dual-domain decoder
To demonstrate the effectiveness of the cross-dual-domain decoder, an experiment was designed to compare the AUROC value and the AUPR value obtained by DFDRNN with those obtained by five variants of DFDRNN, labeled as $s_{r},s_{d},s^{non},s^{non}_{r}$, and $s^{non}_{d}$. Here, $s_{r}$ (Equation 15), $s_{d}$ (Equation 16), and $s^{non}$ represent the models derived from DFDRNN by substituting the cross-dual-domain decoder with a single drug-domain decoder, a single disease-domain decoder and a non-cross dual-domain decoder encompassing decoding results of both the drug and disease domains (as shown in Figure 3), respectively. Similarly, $s^{non}_{r}$ and $s^{non}_{d}$ are models obtained from $s^{non}$ by replacing the non-cross dual-domain decoder with a single drug-domain decoder and a single disease-domain decoder, respectively. The specific implementation of the non-cross dual-domain decoder is described as follows.
${s^{non}_{r}}$ $=\texttt{sigmod}(H^{r}_{a}\times(H^{d}_{a})^{T})$ (22)
${s^{non}_{d}}$ $=\texttt{sigmod}(H^{d}_{s}\times(H^{r}_{s})^{T})$ (23)
${s^{non}}$ $={(s^{non}_{r}+(^{non}_{d})^{T})/2.0}$ (24)
The heatmaps, a crucial visual representation of the results, are shown in Figure 4 across the four datasets: Fdataset, Cdataset, Ldataset, and LRSSL. DFDRNN’s consistent outperformance in achieving the highest AUROC and AUPR values across all four datasets underscores the significance of the cross-dual-domain decoder. Furthermore, the performance of $s_{r}$, $s_{d}$, $s^{non}_{r}$, and $s^{non}_{d}$ consistently lags behind that of DFDRNN and $s^{non}$, further highlighting the effectiveness of dual-domain decoding.
Figure 4: Performance of different decoding methods: (a) AUROC heatmap; (b) AUPR heatmap.
3.7 Identification of candidates for new diseases
In late 2019, the discovery of COVID-19 presented a challenge in promptly researching and developing suitable drugs for treatment. Repurposing already approved drugs, such as remdesivir, favipiravir, and ribavirin, played a significant role in the timely control of the epidemic singh2020drug . Finding suitable drugs for emerging diseases is a complex and crucial task. To evaluate DFDRNN’s predictive ability in identifying potential candidate drugs for new diseases, we conducted leave-one-out cross-validation (LOOCV) on the Fdataset dataset. Specifically, for each disease $d_{x}$, all drug-disease associations related to $d_{x}$ were excluded from the analysis and treated as test samples. The remaining drug-disease associations were used as training samples, relying solely on the similarity between disease $d_{x}$ and other diseases to predict its candidate drugs.
Figure 5: The performance of all methods in predicting potential drugs for new diseases on Fdataset: (a) Receiver operating characteristic (ROC) curves of prediction results obtained by applying DFDRNN and other competitive methods; (b) precision-recall (PR) curves of prediction results obtained by applying DFDRNN and other competitive methods.
In Figure 5, the performance of DFDRNN in identifying potential drugs for new diseases on the Fdataset is illustrated. The experimental results show that DFDRNN achieved the highest AUROC of 0.8269, surpassing the second model NCHDDA by 0.99%. Additionally, DFDRNN had the second-best performance in terms of AUPR, coming in behind only DRWBNCF.
3.8 Case study
To assess the real-world performance of DFDRNN, we conducted case studies on two diseases (Alzheimer’s disease (AD) and Parkinson’s disease (PD)) from the Fdataset. We utilized all drug-disease associations in the Fdataset as the training set to train DFDRNN and predict the association probabilities for all unknown drug-disease pairs. Subsequently, we ranked the drugs in descending order based on the association probabilities and obtained the top 10 predicted drugs for AD and PD. To confirm the accuracy of DFDRNN’s predictions for AD and PD, we adopted highly reliable sources and clinical trials as references, including DB knox2024drugbank , PubChem kim2016pubchem , CTD davis2023CTD , DrugCentral avram2021drugcentral , and ClinicalTrials.
AD, the most prevalent form of progressive dementia in the elderly, is a neurodegenerative disorder. Table 5 displays the top 10 drugs predicted by DFDRNN for AD. Among these, Amantadine is the first drug predicted to hold therapeutic potential for AD. Amantadine, commonly used as an antiviral and antiparkinson drug, exhibits actions as an antiviral, antiparkinson, dopaminergic, and analgesic agent. It can also be utilized for the prevention or symptomatic treatment of influenza A. Previous studies have shown improvements in the mental state of two AD patients following Amantadine treatment erkulwater1989amantadine . Carbamazepine functions as an anticonvulsant and analgesic drug. Earlier studies have suggested that the combined therapy of carbamazepine and haloperidol holds promise in the clinical management of elderly AD patients lemke1995effect . DFDRNN predicts an association with carbamazepine, which has been validated by CTD. Moreover, all the other drugs listed in Table 5 have been confirmed by relevant authoritative data sources, with a success rate of 100%.
Table 5: The top 10 DFDRNN-predicted candidate drugs for AD
[…]
2 DB00564 Carbamazepine CTD
3 DB04844 Tetrabenazine CTD
4 DB00682 Warfarin CTD
5 DB01241 Gemfibrozil CTD/PubChem/ ClinicalTrials
6 DB00160 Alanine CTD/ClinicalTrials
7 DB00903 Etacrynic acid CTD
8 DB00928 Azacitidine CTD
9 DB00482 Celecoxib CTD
10 DB01262 Decitabine CTD
[…]
Table 6: The top 10 DFDRNN-predicted candidate drugs for PD
[…]
2 DB01068 Clonazepam CTD/ClinicalTrials
3 DB01235 Levodopa DB/CTD/PubChem/ DrugCentral/ClinicalTrials/
4 DB00810 Biperiden DB/CTD/PubChem/ DrugCentral
5 DB00424 Hyoscyamine DB
6 DB00575 Clonidine ClinicalTrials
7 DB00915 Amantadine DB/PubChem/DrugCentral/ ClinicalTrials
8 DB00989 Rivastigmine DB/CTD/PubChem/ DrugCentral/ClinicalTrials
9 DB00376 Trihexyphenidyl DB/CTD/PubChem/ DrugCentral
10 DB00215 Citalopram CTD
[…]
Parkinson’s disease (PD) is the second most common neurogenic disorder after AD. Table 6 presents the top 10 potential candidate drugs for PD predicted by DFDRNN. Among these candidates, Haloperidol, an antipsychotic agent, has been confirmed to be associated with PD by DB, CTD, PubChem, and ClinicalTrials data sources. Additionally, Levodopa, a dopamine precursor used in PD management, has also been confirmed by multiple data sources. The fact that the predictions made by DFDRNN for the top 10 PD-related drugs have been validated by authoritative data sources provide reassurance about the reliability of the model.
4 Conclusion
Our study introduces DFDRNN, a neural network model designed for drug repositioning. DFDRNN uses IntraDDFE and InterDDFE modules to extract the similarity and association features of drugs (or diseases), respectively, achieving precise encoding of drugs (or diseases). Compared with the IntraDDFE module, the InterDDFE module changes the similarity (or association) feature into the association (or similarity) feature. In both modules, we use the self-attention mechanism to aggregate the neighbor information of drugs (and diseases). Additionally, the DFDRNN model fuses three similarity (or association) features into a new similarity (or association) feature and uses a layer attention mechanism to calculate the final similarity (or association) feature. Finally, a cross-dual-domain decoder process is adopted to predict drug-disease associations in both the drug and disease domains. DFDRNN performs best in 10-fold cross-validation across four benchmark datasets and achieves the highest AUROC in leave-one-out cross-validation. The experimental results confirm the effectiveness of our proposed DFDRNN model in drug repositioning.
In the future, we will consider incorporating more biological entities, such as miRNA and proteins, for drug repositioning. Additionally, we will explore more accurate methods to calculate drug (or disease) similarity to reduce data noise.
Key Points We propose DFDRNN, a new deep learning method for drug repositioning, which simultaneously extracts two kinds of features: the similarity feature and the association feature to represent drugs and diseases in each network layer. Based on two dual-feature extraction modules: the intra-domain dual-feature extraction (IntraDDFE) module and the inter-domain dual-feature extraction (InterDDFE) module, DFDRNN captures the potential relationships between the features of drugs and diseases. A cross-dual-domain decoder is proposed to implement a cross-computation between two encoded features to redict drug-disease associations. DFDRNN outperformed the state-of-the-art methods on four benchmark datasets.
5 Acknowledgments
This work was supported by the National Key R&D Program of China (No. 2021ZD0112400) and the National Natural Science Foundation of China under grants 62172072 and 62272115.